### v3

Abstract
Nighttime camera-based depth estimation is a highly challenging task, especially for autonomous driving applications, where accurate depth perception is essential for ensuring safe navigation. Models trained on daytime data often fail in the absence of precise but costly LiDAR. Even vision foundation models trained on large amounts of data are unreliable in low-light conditions. In this work, we aim to improve the reliability of perception systems at night time. To this end, we introduce Light Enhanced Depth (LED), a novel, cost-effective approach that significantly improves depth estimation in low-light environments by harnessing a pattern projected by high definition headlights available in modern vehicles. LED leads to significant performance boosts across multiple depth-estimation architectures (encoder-decoder, Adabins, DepthFormer, Depth Anything V2) both on synthetic and real datasets. Furthermore, increased performances beyond illuminated areas reveal a holistic enhancement in scene understanding. Finally, we release the Nighttime Synthetic Drive Dataset, a synthetic and photo-realistic nighttime dataset, which comprises 49,990 comprehensively annotated images. To facilitate further research, both synthetic dataset and code are publicly available at https://simondemoreau.github.io/LED/.
1 Introduction
Figure 1: Light Enhanced Depth (LED) is a novel method that harnesses high-definition headlights’ pattern projected onto the scene to enhance nighttime depth estimation from RGB images. We release a synthetic nighttime dataset with high beam and pattern-illuminated images, along with comprehensive ground truth annotations, to advance research in nighttime perception.
Adverse conditions, such as harsh weather or nighttime, pose significant challenges to many computer vision applications. Despite impressive progress in perception systems for autonomous driving, enabled by powerful deep neural architectures and training techniques, the challenges of nighttime navigation persist. Accurate depth estimation is a crucial aspect of perception, profoundly impacting overall scene perception and comprehension, especially at night [Silberman and Fergus(2011)]. While LiDAR sensors offer high accuracy, their widespread adoption is impeded by their substantial cost. Additionally, in humid weather, LiDAR effectiveness can be significantly degraded due to beam reflection and increased noise levels [Bijelic et al.(2018a)Bijelic, Gruber, and Ritter, Bijelic et al.(2018b)Bijelic, Gruber, and Ritter].
Cameras are powerful, cost-effective sensors that capture rich information about the environment and are deployed in most modern vehicles. Camera-based perception solutions display high reliability and accuracy [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Bian et al.(2019)Bian, Li, Wang, Zhan, Shen, Cheng, and Reid, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Godard et al.(2017)Godard, Mac Aodha, and Brostow, Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023), Li et al.(2023)Li, Chen, Liu, and Jiang], but mostly on clear daylight conditions: they struggle under distribution shift and low-light conditions. While foundation models [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Bochkovskii et al.(2024)Bochkovskii, Delaunoy, Germain, Santos, Zhou, Richter, and Koltun, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] show great improvement for depth estimation on multiple domains, they are still not robust to nighttime images that are long-tail in the training data distribution. The availability of labeled data necessary for training high-capacity deep neural networks is an additional challenge. For the task of depth estimation, some methods rely on supervised learning [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023), Li et al.(2023)Li, Chen, Liu, and Jiang, Lee et al.(2019)Lee, Han, Ko, and Suh], but most use self-supervised methods [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Poggi et al.(2018)Poggi, Aleotti, Tosi, and Mattoccia, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Yin and Shi(2018), Shu et al.(2020)Shu, Yu, Duan, and Yang, Bian et al.(2019)Bian, Li, Wang, Zhan, Shen, Cheng, and Reid, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al.] since obtaining ground truth depth information is expensive. While dedicated approaches for specific conditions have emerged in recent years [Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al., Vankadari et al.(2020)Vankadari, Garg, Majumder, Kumar, and Behera, Spencer et al.(2020)Spencer, Bowden, and Hadfield, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang], they still only partially mitigate these issues. To the best of our knowledge, they remain limited to self-supervised strategies, due to the lack of publicly available, large-scale annotated datasets suitable for nighttime depth estimation. To address this gap, this paper releases a synthetic nighttime dataset annotated with dense depth maps, along with additional labels.
In this paper, we introduce Light Enhanced Depth (LED), a novel approach that significantly improves depth estimation in low-light environments, ensuring enhanced accuracy and reliability for autonomous vehicles. High-Definition (HD) headlights, commonly found in modern vehicles, have shown promising results in scene perception research [Waldner et al.(2019)Waldner, Kramer, and Bertram, Waldner et al.(2020)Waldner, Kramer, and Bertram, Waldner and Bertram(2021)]. Drawing inspiration from active stereovision [Fanello et al.(2016)Fanello, Rhemann, Tankovich, Kowdle, Escolano, Kim, and Izadi, Riegler et al.(2019)Riegler, Liao, Donne, Koltun, and Geiger, Li et al.(2022)Li, Monno, and Okutomi, Baek and Heide(2021)], we harness HD headlights to project a pattern into the scene, guiding the network and thereby improving performance.
Our contributions can be summarized as follows:
$\bullet$ Architecture-agnostic enhancement: LED can be applied to any depth estimation architecture to improve nighttime performances (RMSE: -11% on encoder-decoder, -24.06% on Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka], -8.00% on DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang], -15.5% on Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]).
$\bullet$ Data-efficiency: LED-trained models outperform the others with only 20% of training data.
$\bullet$ Real prototype: LED demonstrates great performances improvements on our in-house dataset, collected using a real car-mounted prototype.
$\bullet$ Dataset: We provide the Nighttime Synthetic Drive Dataset, a photorealistic synthetic dataset, comprising 49,995 comprehensively annotated nighttime images to foster future investigations in nighttime perception.
2 Related Work
2.1 Depth Estimation from Camera
Supervised Learning. Recent advances in supervised depth prediction leverage transformer architectures and attention mechanisms [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023), Li et al.(2023)Li, Chen, Liu, and Jiang, Lee et al.(2019)Lee, Han, Ko, and Suh] to improve performances. The use of bins [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Agarwal and Arora(2023)] improves depth prediction accuracy even with limited datasets available [Geiger et al.(2012)Geiger, Lenz, and Urtasun, Caesar et al.(2020)Caesar, Bankiti, Lang, Vora, Liong, Xu, Krishnan, Pan, Baldan, and Beijbom, Maddern et al.(2017)Maddern, Pascoe, Linegar, and Newman, Nathan Silberman and Fergus(2012)] compared to self-supervised methods [Godard et al.(2019)Godard, Mac Aodha, Firman, and Brostow, Poggi et al.(2018)Poggi, Aleotti, Tosi, and Mattoccia, Zhou et al.(2017)Zhou, Brown, Snavely, and Lowe, Yin and Shi(2018), Bian et al.(2019)Bian, Li, Wang, Zhan, Shen, Cheng, and Reid, Guizilini et al.(2020)Guizilini, Ambrus, Pillai, Raventos, and Gaidon, Shu et al.(2020)Shu, Yu, Duan, and Yang]. However, these methods are confined to daytime scenarios. Nighttime depth estimation approaches [Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al.] rely on self-supervised methods, due to the lack of large, annotated nighttime datasets. In response, we release the Nighttime Synthetic Drive Dataset to support the development of nighttime-specific methods.
Vision Foundation Models. Depth foundation models [Yang et al.(2024a)Yang, Kang, Huang, Xu, Feng, and Zhao, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao, Bochkovskii et al.(2024)Bochkovskii, Delaunoy, Germain, Santos, Zhou, Richter, and Koltun, Ke et al.(2024)Ke, Obukhov, Huang, Metzger, Daudt, and Schindler] have achieved impressive generalization by leveraging extensive synthetic and real-world training data. However, we show in section˜5.4 that these models perform poorly in zero-shot evaluations on our nighttime dataset. It highlights the need for models specialized in nighttime conditions and underscores the benefits of our method.
Nighttime Domain Adaptation. Nighttime depth estimation can be approached as a domain adaptation problem [Gasperini et al.(2023)Gasperini, Morbitzer, Jung, Navab, and Tombari], focusing on aligning features between daytime and nighttime images [Vankadari et al.(2020)Vankadari, Garg, Majumder, Kumar, and Behera, Spencer et al.(2020)Spencer, Bowden, and Hadfield, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang].  Our method specifically adapts to nighttime scenes features by leveraging the informative deformations of the light pattern emitted by a vehicle’s HD headlamp.
2.2 Light for Perception
Active Stereovision. Depth estimation based on active stereovision involves disparity measurement. Unlike traditional stereovision, disparity is computed between an image and a pattern projected onto the scene. Recent methods [Fanello et al.(2016)Fanello, Rhemann, Tankovich, Kowdle, Escolano, Kim, and Izadi, Riegler et al.(2019)Riegler, Liao, Donne, Koltun, and Geiger, Li et al.(2022)Li, Monno, and Okutomi, Baek and Heide(2021)] use deep learning models designed to take patterns and images as input. While active stereovision offers valuable scale information, it has demonstrated significant performance degradation when deployed outdoors, due to high ambient lighting and the low power of projectors [Mertz et al.(2012)Mertz, Koppal, Sia, and Narasimhan, Gupta et al.(2013)Gupta, Yin, and Nayar, Brahmanage and Leung(2019)]. We avoid this thanks to our nighttime environment and the utilization of high-powered HD headlights.
HD Lighting. A series of works [De Charette et al.(2012)De Charette, Tamburo, Barnum, Rowe, Kanade, and Narasimhan, Tamburo et al.(2014)Tamburo, Nurvitadhi, Chugh, Chen, Rowe, Kanade, and Narasimhan] highlight HD lighting potential applications and particularly its ability to design anti-glare systems while optimizing illumination for enhanced driver visibility, even in adverse conditions such as rain or snow. More recently, [Waldner et al.(2022)Waldner, Müller, and Bertram] propose reducing overall scene illumination to decrease power consumption while maintaining object detection capabilities. [Waldner et al.(2019)Waldner, Kramer, and Bertram, Waldner et al.(2020)Waldner, Kramer, and Bertram, Waldner and Bertram(2021)] also have proposed hardware-in-the-loop simulation for HD headlights. We introduce a novel application of HD headlights to improve understanding of overall scene geometry and enhance depth estimation. We achieve this by projecting a pattern onto the nighttime scene, providing guidance for the model.
3 Method
We improve depth estimation from monocular camera at nighttime by leveraging a pattern projected in front of the vehicle with HD headlights.
3.1 HD Pattern and Headlights
Networks designed for active stereovision exploit the disparity between the projected pattern and the camera’s view to estimate distances within the scene. With LED, a model identifies pattern’s areas that deviate from the implicitly learned reference pattern.
Figure 2: Pattern deformation in the scene: Left: For explanatory purpose, we project the pattern on a depth image. It demonstrates trapezoidal deformations on horizontal surfaces and undistorted squares on vertical surfaces. Complex deformations on the car reveal insights into its geometry. Right: Pattern projected on a wall at 10 m (darker) and 100 m (lighter). Square sizes increases with distance.
Unlike traditional active stereovision, which relies on infrared lasers to project points, we harness an HD headlight. Similar to a projector, this headlight can dynamically project any image or shape onto the scene. To prevent pattern overlap, we use only the left headlight.
We employ a regular checkerboard pattern, because of its high contrast and sharp discontinuities. The dense concentration of corners and transitions makes it easily detectable by convolutional neural networks [Baker et al.(2020)Baker, Lu, Erlikhman, and Kellman]. Upon projection, the checkerboard is deformed according to surface shapes (from the camera’s perspective). Horizontal planes stretch the pattern into a trapezoid, while vertical surfaces parallel to the image plane cause no deformation (see fig.˜2). Non-planar surfaces, such as cars, result in more complex distortions. These deformations provide crucial geometric cues to the network, improving depth estimation accuracy. Additionally, the emitted light is not colimated: as the projection extends further, the light becomes more divergent, causing pixels to appear larger (see fig.˜2), thus providing useful depth cues to the network.
3.2 Architectures
Our method is architecture-agnostic. By relying on a single pattern, the model can implicitly learn it, eliminating the need for a dedicated network architecture. To demonstrate the benefits of projecting an HD pattern for depth estimation, we leverage a simple encoder-decoder model [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox]. In spite of its straight-forward architecture, it achieves results on par with or superior to other state-of-the-art (SOTA) methods (see section˜5.1). We also successfully apply our method to more complex SOTA architectures, such as DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang] and Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka], with similar performance gains (see section˜5.1). Finally, we show that finetuning Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] with LED, on a reduced amount of training samples, enable robust and precise depth estimation.
3.3 Implementation Details
We implement the encoder-decoder in PyTorch [Paszke et al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan, Killeen, Lin, Gimelshein, Antiga, Desmaison, Kopf, Yang, DeVito, Raison, Tejani, Chilamkurthy, Steiner, Fang, Bai, and Chintala]. The model is trained from scratch during 70 epochs using the AdamW optimizer [Loshchilov and Hutter(2018)], with a batch size of 32 and a learning rate of $10^{-3}$. Input and output resolutions are set to $\times320$ px. Similar to [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani], our learning objective combines losses based on Log L1, gradient and normals. More information is available in the supplementary. The selection of the best epoch is based on the Root Mean Square Error (RMSE). The training process takes 4 hours on a single Nvidia RTX 4090.
We implement Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka] using the official code and rely on the toolbox [Li(2022)] for DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang]. We apply Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] official fine-tuning code to adapt the model for metric depth estimation. The only changes are limited to dataset format compatibility. The training setup follows the original papers’ recommendations, excluding data augmentation to preserve the HD pattern from cropping.
Figure 3: Nighttime Synthetic Drive Dataset examples: (A) depicts HD pattern and (B) high-beam illumination. Ground truth annotations include dense depth maps, semantic segmentation, instance segmentation labels and bounding boxes.
4 Dataset
4.1 Nighttime Synthetic Drive Dataset (NSDD)
[…]
We choose High Beam (HB) as the comparison baseline, considering it the maximum normal lighting condition. Since high beam illumination exceeds that of the checkerboard pattern, performance improvements are attributed solely to the pattern’s contribution. In each frame, vehicles, pedestrians, and traffic signs are randomized. To simulate light interference, we include other car headlights and randomize ambient light levels between 0 and 10 lux.
Figure 4: Simulation of HD pattern: Top-left: control matrix of the HD headlight. Bottom-left: photometry considering aberrations created by the headlight lens. Right: Resulting image using the photometry. The area outlined in red is the region of interest.
The dataset consists of 5 different maps (3 train / 1 val / 1 test), with 4,999 frames generated for both pattern and HB illuminated images in each. These maps represent major cities. Both parts of the dataset share the same randomization code, making the domains comparable.
[…]
4.2 Real-world Dataset
We validate LED on a real-world, in-house dataset collected using a car-mounted prototype. It comprises 50,000 images (70% train / 15% val / 15% test) from populated urban and rural roads, evenly split between Low Beam (LB) illumination and checkerboard pattern with 0.5°, 0.25° and 0.125° cells’ dimension. Ground truth is obtained from LiDAR data using the DOC-Depth method [de Moreau et al.(2025)de Moreau, Mathias, Hassan, Yasser, Andrei, Hafid, Fabien, et al.] and Exwayz software [Exwayz(2024)]. More details are available in supplementary materials.
5 Experiments
Figure 5: Qualitative results on NSDD. LED results exhibit higher precision, and more accurate object boundaries and shapes compared to HB. Red boxes indicate enhanced regions.
Pattern RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
[…]
HB 6.1204 0.0903 0.0233 0.1353 13.4847 3.3579 0.9489 0.9812 0.9908
LED 5.4259 0.1996 0.0188 0.1253 12.4900 16.1224 0.9603 0.9846 0.9927
[…]
HB 7.3520 0.1360 0.0320 0.1230 10.1880 4.2770 0.8790 0.9440 0.9740
LED 5.5830 0.0690 0.0240 0.0920 7.7290 2.5090 0.9210 0.9680 0.9840
[…]
HB 5.8528 0.0687 0.0281 0.1205 11.3841 0.8328 0.9432 0.9802 0.9913
LED 5.3845 0.1215 0.0276 0.1111 10.4146 3.4015 0.9497 0.9838 0.9950
Table 1: Comparison of depth estimation performances on NSDD: HB models are trained on high beam data and LED on HD light pattern. LED models outperform HB models across all metrics, exceptions are for Abs Rel and Sq Rel. Metrics are computed in the ROI.
We evaluate our method through extensive experiments on the Nighttime Synthetic Drive Dataset. We validate the contribution of our light pattern in boosting performance, both inside and outside illuminated area. We apply our method to the encoder-decoder and other SOTA approaches: Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka], DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang] and Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao], showing its agnosticity. Finally, we evaluate LED robustness beyond its training domain and its capabilities on real-world scenarios. Metrics used align with prior works [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Li et al.(2023)Li, Chen, Liu, and Jiang, Spencer et al.(2020)Spencer, Bowden, and Hadfield, Liu et al.(2021)Liu, Song, Wang, Liu, and Zhang, Wang et al.(2021)Wang, Zhang, Yan, Li, Xu, Li, and Yang, Zheng et al.(2023)Zheng, Zhong, Li, Gao, Zheng, Jin, Wang, Zhao, Zhou, Zhang, et al.].
Model RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
(HB) O-ROI 9.2533 0.1598 0.0249 0.2047 20.4447 11.4068 0.9385 0.9714 0.9828
(LED) O-ROI 9.0070 0.1295 0.0221 0.2027 20.2348 8.7407 0.9427 0.9723 0.9828
(HB) Full 8.9702 0.1521 0.0247 0.1988 19.8572 10.5225 0.9396 0.9725 0.9836
(LED) Full 8.6963 0.1371 0.0217 0.1965 19.6137 9.5517 0.9446 0.9737 0.9839
Table 2: Encoder-decoder performance beyond ROI on NSDD: O-ROI stands for Outside ROI, where the evaluation mask is the inverse of the ROI. For Full, metrics are computed on the entire image. (LED) models trained on HD pattern outperform (HB) ones in all metrics.
Figure 6: Performance comparison of LED and HB in various settings and metrics on NSDD. (a) Performance over distance; (b) Performance over training set size; (c) Robustness across domains: The amount of training data is fixed and composed of a given ratio of LED and HB data. Overall, LED achieves better long-range results while being more data-efficient. It also demonstrates improved performance across domains.
5.1 HD Pattern Impact
We assess LED impact on depth estimation by comparing results with and without pattern. As reference, we take models trained on high beam (HB), while the variants use our HD pattern (LED). Quantitative and qualitative results are presented in table˜1 and fig.˜5.
Quantitative Results. We evaluate our method on the encoder-decoder and two SOTA architectures commonly used in daytime scenarios, Adabins [Bhat et al.(2021)Bhat, Alhashim, and Wonka] and DepthFormer [Li et al.(2023)Li, Chen, Liu, and Jiang]. Their use of transformer architectures offers valuable insights into LED behavior on more recent architectures compared to the convolutional ones that are common in embedded devices.
Comparing the encoder-decoder LED with HB (table˜1), we observe a significant improvement in metrics within the region-of-interest (ROI): LED yields -11% RMSE. Using LED, the model provides more accurate distance estimations. However, its errors are often related to pattern occlusions, which affect both near and far objects, resulting in a degradation of relative metrics (Abs Rel, Sq Rel). All metrics assessing global precision are improved.
Compared to the SOTA approaches [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Li et al.(2023)Li, Chen, Liu, and Jiang], LED-trained encoder-decoder either outperforms or matches the performance of their HB models. Given that the encoder-decoder is less tailored for depth estimation than more intricate SOTA designs, this result underscores the substantial improvements yield by LED, demonstrating that using the HD informative pattern even with a straightforward architecture is promising for challenging nighttime scenes.
With LED, Adabins demonstrates a substantial improvement compared to the HB model: -24.06% RMSE and -6.70% Abs Rel. Conversely, DepthFormer showcases -8.00% in RMSE but +5.28% in Abs Rel. Note that the most significant enhancement is observed in Adabins, which have the worst results by night on HB.
Improvements seen across diverse architectures confirm LED agnosticity to architectures. It implies potential effectiveness with future methods.
Performance Over Distance. Limited nighttime visibility makes depth estimation of distant objects challenging. In fig.˜6 (a), we show the performance of the encoder-decoder over distances. LED maintains performance for close object and exhibits greater enhancement at longer ranges, thus addressing depth estimation of distant objects at night.
[…]
LED 5.497 0.209 12.635
HB 6.443 0.095 13.723
HL 6.668 0.068 14.308
VL 7.360 0.172 15.774
Table 3: Performances of the encoder-decoder using various patterns on NSDD. VL and HL refers to vertical and horizontal lines. Results are obtained from 50% of NSDD training data.
Pattern Study. We generated additional data with various illumination patterns to compare the performance of the LED checkerboard against other common structured light patterns (horizontal and vertical lines). Table˜3 shows that the checkerboard pattern outperforms the others across most metrics. While horizontal lines show slight improvement in relative metrics, they cause a significant decline in others.
Domain RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
[…]
Table 4: Encoder-decoder performance across domains of NSDD. We denote (training domain $\rightarrow$ testing domain). The model trained on (HB) fails entirely when tested on (LED). Yet, (LED)-trained models are able to estimate depth when tested on the (HB) domain.
Pattern # Images RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
HB Zero-shot 19.058 0.430 0.243 0.587 27.004 8.134 0.086 0.186 0.785
LED 18.023 0.420 0.232 0.563 25.500 7.779 0.096 0.214 0.867
HB 100 6.499 0.075 0.033 0.133 12.894 0.992 0.929 0.977 0.991
LED 6.198 0.073 0.032 0.131 12.182 0.898 0.933 0.977 0.991
HB 500 5.447 0.058 0.025 0.112 10.629 0.688 0.950 0.985 0.995
LED 5.225 0.057 0.025 0.108 10.038 0.648 0.953 0.986 0.995
HB 1000 5.124 0.058 0.026 0.107 10.038 0.607 0.953 0.985 0.995
LED 4.328 0.044 0.020 0.089 8.374 0.430 0.962 0.991 0.997
Table 5: Depth Anything V2 performances on NSDD. The model fails when used as zero-shot, but it correctly estimates depth with few-shot learning. LED always outperforms HB.
5.2 Improving Global Scene Understanding
We investigate the impact of LED on depth estimation beyond the ROI, highlighting its contribution to a more holistic scene understanding. Results are presented in table˜2.
Examining depth estimation both outside the ROI and across the entire image, we observe a consistent performance improvement: -2.66% RMSE outside the ROI and -3.05% RMSE on the entire image. We also report enhancement in relative metrics: Abs Rel (resp. -3.03% and -1.50%) and Sq Rel (resp. -23.27% and -9.23%). These results suggest that the HD headlight pattern provides valuable information, such as object size and scale, leading to enhanced overall scene understanding and more accurate depth estimation.
5.3 Annotation-efficient Learning
Collecting annotated nighttime data is challenging and costly. We explore whether pattern extra guidance can reduce the need for training data (refer to fig.˜6 (b)). Our results show that encoder-decoder model trained with less than 20% of the LED data outperforms one trained on the full HB dataset, reaching near-peak performance with just 50%. In addition, we observe that incorporating just 10% of LED data into the HB training set enable the network to learn relevant features and enhance its performances when used with pattern (see fig.˜6 (c)). Therefore, any vehicle equipped with HD headlights can apply LED by adding only a few pattern images in their training set, thereby reducing the cost of specialized data.
5.4 Few-shot Learning
Foundation models show strong zero-shot performance across various cameras and content types. However, when tested on NSDD, Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] performs poorly, with an RMSE of $\sim$19 m (see table˜5), highlighting the complexity of nighttime data and the need for specialized models. It performs slightly better on LED images, possibly due to its ability to leverage the pattern’s geometric cues. Fine-tuning with as little as 100 images is sufficient for adapting the network to our distribution and produce consistent depth predictions on both LED and HB. LED-models demonstrate significant improvements across all metrics (-15.5% RMSE with 1,000 images). This underlines the effectiveness of our method in enhancing nighttime depth estimation even on foundation models. Results on our real-world dataset are provided in the supplementary material.
Pattern RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
[…]
LB 14.199 0.187 0.072 0.292 28.938 5.620 0.772 0.909 0.956
LED 0.5° 11.089 0.119 0.050 0.213 21.251 2.304 0.864 0.950 0.977
LED 0.25° 8.695 0.109 0.040 0.190 18.635 2.128 0.899 0.963 0.984
LED 0.125° 10.154 0.096 0.040 0.190 18.862 2.132 0.907 0.969 0.984
[…]
LB 8.777 0.138 0.050 0.186 17.170 3.365 0.864 0.957 0.983
LED 0.5° 6.810 0.101 0.041 0.152 14.380 1.376 0.890 0.969 0.989
LED 0.25° 5.621 0.082 0.030 0.126 11.735 1.163 0.923 0.978 0.992
LED 0.125° 5.727 0.076 0.030 0.116 10.754 1.261 0.940 0.987 0.995
[…]
LB 7.637 0.104 0.044 0.160 14.617 1.401 0.891 0.972 0.990
LED 0.5° 7.017 0.109 0.047 0.161 14.044 1.260 0.877 0.9697 0.989
LED 0.25° 6.304 0.102 0.044 0.146 12.300 1.079 0.885 0.973 0.993
LED 0.125° 5.785 0.082 0.037 0.126 10.673 0.840 0.921 0.981 0.992
Table 6: Performance comparison on real-world data using various pattern resolutions. LED models outperform LB on all metrics, showing LED suitability on real-world scenarios.
5.5 Robustness Across Domains
[…]
5.6 Real-world Scenarios
[…]
To account for objects passing through the pattern illumination, we investigate the impact of smaller checkerboard cells. Using a size of 0.25°, we observe great enhancement across all metrics (-17.4% RMSE, -18.8% Abs Rel against 0.5°). Further increasing the resolution to 0.125° does not lead to significant improvements.
[…]
In contrast, Depth Anything V2 achieves its best performance with a 0.125° resolution, as it is better suited to capturing fine details. When fine-tuned with only 1,000 images, we observe -24.3% RMSE compared to the LB baseline. Performances shown by DepthFormer and Depth Anything V2 underpin the benefits of LED in complex real-world scenarios.
6 Limitations and Future Works
The method’s reliance on a single reference pattern ensures agnosticism but limits its flexibility. Designing specialized architectures that treat the pattern as an input could enable dynamic pattern optimization and improve generalization.
We tested varying projector-camera distances (real: 70 cm, NSDD: 150 cm), showing LED robustness on similar car models. Future work will assess its usage with larger variations.
LED patterns may cause glare for other road users. Leveraging [Tamburo et al.(2014)Tamburo, Nurvitadhi, Chugh, Chen, Rowe, Kanade, and Narasimhan] to mask them could mitigate this. Future work should assess both safety concerns and the impact on performance.
7 Conclusion
We introduce LED, a method for enhancing nighttime depth estimation by leveraging high-definition light patterns projected by modern vehicle headlights. Through extensive experiments, we demonstrate significant improvements in depth perception, both within and beyond illuminated areas. Our method’s versatility is highlighted by its successful integration with two state-of-the-art architectures, Adabins and DepthFormer, as well as the foundation model Depth Anything V2. Moreover, LED shows promising real-world performance. We also release the Nighttime Synthetic Drive Dataset, comprising 49,995 fully annotated images. We hope it will serve as a valuable resource for the research community, supporting exploration of various nighttime perception tasks.
Acknowledgment
We would like to thank Amandine Brunetto for her help and insights. We also thank Naimeric Villafruela from the Fablab of Mines Paris for his assistance in creating the real-world prototype. In addition, we extend our warm thanks to the Exwayz team for the excellent collaboration on the real-world ground truth processing. This work was granted access to the HPC resources of IDRIS under the allocation 2025-AD011015334 made by GENCI.
Supplementary Material
[…]
Mines Paris - PSL University
[…]
Paris, France
LED: Light Enhanced Depth Estimation at Night
A Safety and Regulation
While using readily available hardware, LED is a research project not meant to be deployed on cars right away. Future works should assess potential safety issues. To maximize their safety, autonomous vehicles needs multiple perception mechanisms that ensure redundancy, compensating for blind spots of other sensors on the car, e.g., LiDAR in rain conditions. One can imagine that this light pattern could be turned off in crowded areas, or when detecting an incoming car after it was initially detected at a longer distance, ensuring safety. Regarding regulation, HD headlight technology is novel and the European regulation has authorized just recently the projection of specific HD pattern onto the road. Thus, laws are moving in this direction and evolving with the technology.
B Encoder-Decoder Details
B.1 Encoder-Decoder Architecture
LED uses a single pattern implicitly learned by the model, making it architecture-agnostic. This characteristic is demonstrated in the main paper (see section section˜5.1) by applying LED to multiple state-of-the-art architectures [Bhat et al.(2021)Bhat, Alhashim, and Wonka, Li et al.(2023)Li, Chen, Liu, and Jiang, Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao]. To prove our concept and conduct experiments, we opt for an encoder-decoder architecture with skip connections [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox]. The detailed architecture is shown in fig.˜S1. Despite the simplicity of this architecture, we show that LED enables the network to achieve performances comparable to or better than other tested SOTA architectures trained without our method.
Figure S1: Detailed encoder-decoder architecture.
B.2 Learning Objectives
To train the encoder-decoder, we adopt a combination of loss functions inspired by [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani]. The primary loss, $\mathcal{L}_{depth}$, is the Log L1 loss presented in eq.˜1. It measures the error between the estimated depth, $d_{i}$, and the corresponding ground truth, $g_{i}$. This variant of L1 loss attributes less significance to errors occurring at greater distances. This adjustment aligns with the expectation that given errors, e.g., 1 m, should have greater impact when within a few meters of the camera but are more tolerable at extended distances:
$\mathcal{L}_{depth}=\frac{1}{N}\sum_{i=1}^{N}|log(d_{i})-log(g_{i})|.$ (1)
We address edges fidelity, particularly challenging in low-light conditions. To this end, we incorporate a loss, $\mathcal{L}_{grad}$, that specifically emphasizes on gradients. As we believe edges sharpness is important regardless of the distance, we employ a standard L1 Loss expressed in eq.˜2 instead of a logarithmic one used in [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani]. Moreover, our experiments revealed no performance improvements with the Log version. $\nabla_{x}(d_{i})$ and $\nabla_{y}(d_{i})$ respectively represent the spatial derivative of $d_{i}$ along the x and y-axis,
$\mathcal{L}_{grad}=\frac{1}{N}\sum_{i=1}^{N}|\nabla_{x}(d_{i})-\nabla_{x}(g_{i})|+|\nabla_{y}(d_{i})-\nabla_{y}(g_{i})|.$ (2)
Similar to [Hu et al.(2019)Hu, Ozay, Zhang, and Okatani], we want to ensure accurate surfaces representation in depth maps.
The depth normals are estimated at each pixel using $n_{i}^{a}\equiv[-\nabla_{x}(a_{i}),-\nabla_{y}(a_{i}),1]^{\intercal}$. The cosine similarity loss, expressed in eq.˜3, is then employed to compare estimated and ground truth normals. $\langle.,.\rangle$ denotes vector inner product operation:
$\mathcal{L}_{normal}=\frac{1}{N}\sum_{i=1}^{N}|1-\frac{\langle n_{i}^{d},n_{i}^{g}\rangle}{\sqrt{n_{i}^{d},n_{i}^{d}}\sqrt{n_{i}^{g},n_{i}^{g}}}|.$ (3)
Finally, our learning objective can be expressed as eq.˜4, we set $\lambda_{1}=1$ and $\lambda_{2}=1$:
$\mathcal{L}=\mathcal{L}_{depth}+\lambda_{1}\mathcal{L}_{grad}+\lambda_{2}\mathcal{L}_{normal}.$ (4)
B.3 Loss Study
To determine the impact of each individual loss, we conduct an ablation study (see table˜S1). Both L1 and Log L1 losses are tested for $\mathcal{L}_{depth}$ and $\mathcal{L}_{grad}$. Our findings indicate that each selected loss positively influences performance. The combination outlined in section˜B.2 demonstrates the most favorable trade-off between metrics.
$\mathcal{L}_{depth}$ $\mathcal{L}_{grad}$ $\mathcal{L}_{normal}$ Metrics
L1 Log L1 L1 Log L1 RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
5.979 0.191 0.020 0.132 13.032 14.352 0.957 0.982 0.991
5.852 0.187 0.021 0.132 13.122 12.984 0.957 0.983 0.992
5.760 0.201 0.020 0.137 13.616 16.174 0.958 0.983 0.992
5.386 0.202 0.021 0.124 12.373 15.901 0.958 0.985 0.993
5.309 0.206 0.022 0.149 14.862 16.048 0.954 0.981 0.990
5.809 0.186 0.018 0.125 12.210 14.683 0.961 0.985 0.993
5.422 0.200 0.018 0.125 12.503 16.120 0.961 0.985 0.993
Table S1: Loss study. Performance comparison using several combination of losses: 1st best, 2nd best, 3rd best.
C Depth Anything V2 Details
To fine-tune Depth Anything V2 [Yang et al.(2024b)Yang, Kang, Huang, Zhao, Xu, Feng, and Zhao] on our datasets, we used the metric model pre-trained on KITTI. We follow their indications for specializing the model to metric depth estimation. For each given number of training images, the split was selected randomly in the whole dataset. Figure˜S2 illustrates the RMSE of Depth Anything V2 across different numbers of fine-tuning images. We observe that the model is near its performance peak with just 1,000 images. All LED-enhanced models outperform the LB, also this architecture benefits from finer resolution in the pattern, demonstrating its best performances with LED 0.125°.
Figure S2: Few-shot learning performances of Depth Anything V2 on our real dataset. All LED patterns demonstrate superior performances over LB. The model is near its maximal performances with only 1,000 training images.
D ROI Details
D.1 ROI Definition
To assess the impact of the LED pattern, we define a Region of Interest (ROI) representing the illuminated area in most images and calculate our metrics within this region. We center-crop the image to $640\times 640$ px and then resize it to $320\times 320$ px. In the resulting image, the ROI consists of pixels with coordinates satisfying: $20\leq p_{x}\leq 270$ and $165\leq p_{y}\leq 210$, where $p_{x}$ and $p_{y}$ are the pixel coordinates along the $x$ and $y$ axes, respectively.
D.2 ROI-Only Training
To assess the impact of our method, we focus on the illuminated area. Thus, it is reasonable to evaluate performance when trained exclusively within the ROI. We report results of this experiment in table˜S2. We observe that training solely within the ROI enhances Abs Rel and Sq Rel metrics, although other metrics show a decline. Since the network’s ability to estimate depth beyond the ROI is valuable for many applications, we do not pursue this approach of ROI-only training.
Image input RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
Reference 5.4259 0.1996 0.0188 0.1253 12.4900 16.1224 0.9603 0.9846 0.9927
ROI-Only 5.6931 0.1145 0.0199 0.1428 14.2131 4.4541 0.9585 0.9841 0.9932
Table S2: Encoder-decoder performances on NSDD between ROI-only training and reference from section section˜5.1.
Resolution RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
200 px 5.9413 0.1073 0.0206 0.1270 12.6227 3.9736 0.9545 0.9228 0.9921
320 px 5.4259 0.1996 0.0188 0.1253 12.4900 16.1224 0.9603 0.9846 0.9927
640 px 5.7050 0.2109 0.0185 0.1420 14.1092 17.4085 0.9668 0.9837 0.9893
Table S3: Encoder-decoder performances on NSDD at various resolutions.
Figure S3: Encoder-decoder performances on NSDD using multiple resolutions
Figure S4: Example of illuminations pattern tested: high-beam (top-left), checkerboard (top-right), horizontal lines (bottom-left) and vertical lines (bottom-right).
E Resolution Impact
Resolution and performance are usually highly correlated in computer vision, particularly in low-light conditions. To better understand the impact of resolution on our method, we train the encoder-decoder with center-cropped area resized at various resolutions. Results in table˜S3 indicate that increasing resolution up to $640\times 640$ px does not improve performance. Conversely, decreasing resolution to $200\times 200$ px appears to enhance relative metrics while degrading others. Since most valuable cues in our method come from the pattern, these findings suggest that a resolution of $320\times 320$ px offers a favorable trade-off for pattern visibility and scene interpretation within our setup (as shown in fig.˜S3).
Pattern RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
LED 5.5179 0.1977 0.0196 0.1233 12.2798 15.6867 0.9593 0.9849 0.9931
HB 6.0937 0.0867 0.0216 0.1298 12.8895 3.1352 0.9521 0.9824 0.9917
VL 7.3598 0.1723 0.0264 0.1600 15.7743 6.5203 0.9393 0.9751 0.9876
HL 6.6679 0.0681 0.0254 0.1445 14.3077 1.2319 0.9370 0.9776 0.9903
Table S4: Comparison of encoder-decoder performances using various patterns: LED use the checkerboard, HB stands for high-beam, VL and HL are vertical and horizontal lines respectively.
F Examples of Other Patterns
To find a suitable pattern for our application, we generate half-sized datasets featuring commonly used structured light patterns: checkerboard, horizontal, and vertical lines (see fig.˜S4). Comprehensive metrics from this experiment are available in table˜S4, the checkerboard pattern demonstrates significantly superior performances.
G LED Impact On Other Tasks
We introduce the LED lighting pattern for improving performance on geometric tasks. We study its impact on the performance of other tasks that might be running in the same time on the vehicle, taking the example of semantic segmentation. To this end, we train Mask2Former [Cheng et al.(2021)Cheng, Misra, Schwing, Kirillov, and Girdhar] using the available annotation in the Nighttime Synthetic Drive Dataset. Results are reported in table˜S5. LED doesn’t improve, nor degrade performances. We even note a better stability with less variance over runs. Therefore, the similar performance on both domains suggests that LED enhances geometric tasks with limited impact on semantic ones.
Pattern mIoU $\uparrow$
[…]
LB 51.45 $\pm$ 4,00
LED 51.49 $\pm$ 0,76
Table S5: Mask2Former [Cheng et al.(2021)Cheng, Misra, Schwing, Kirillov, and Girdhar] Semantic segmentation results on NSDD.
Test set RMSE $\downarrow$ Abs Rel $\downarrow$ $\text{Log}_{10}$ $\downarrow$ RMSE Log $\downarrow$ SILog $\downarrow$ Sq Rel $\downarrow$ $\delta^{1}$ $\uparrow$ $\delta^{2}$ $\uparrow$ $\delta^{3}$ $\uparrow$
LED (0.5°) 6.859 0.102 0.041 0.153 14.162 1.511 0.894 0.971 0.989
LED (0.25°) 5.787 0.082 0.033 0.126 11.875 1.0839 0.929 0.983 0.994
LED (0.125°) 5.730 0.077 0.030 0.116 10.913 1.149 0.958 0.989 0.995
Table S6: DepthFormer results on real dataset when trained on 33% of each checkerboard resolutions (0.5°, 0.25°, 0.125°).
H Using Multiple Checkerboard Resolutions
Any HD headlight can project LED’s checkerboard, although with a resolution adjustment. To evaluate its impact on model performances, we propose to train a model with 3 checkerboard resolutions (0.5°, 0.25° and 0.125°) evenly split, and evaluate on each one. Table˜S6 demonstrates that the model is able to perform well on all resolutions although with a minor performance degradation compared to single resolution training. This shows our model’s ability to generalize over multiple resolutions, thus, various HD headlights or car models.
Figure S5: Qualitative results on real-world scenarios. LED-trained models are more accurate and object edges are better defined.
I Real World Scenarios
To collect our real-world dataset, we made a prototype on a real car. We used an IDS U3-36L0XC camera and an Ouster OS1-128 LiDAR (Rev 7). Regarding the HD headlight, we opt for the Digital Micromirror Device technology, which offer the greatest resolution (<<0.015°) with a horizontal FOV of 14° and vertical FOV of 7°. All the hardware was mounted on the roof of the car. This novel setup allowed us to test LED under another projector-camera configuration. Ground truth annotation was made using LiDAR data, aggregated and densified following DOC-Depth method [de Moreau et al.(2025)de Moreau, Mathias, Hassan, Yasser, Andrei, Hafid, Fabien, et al.] and thanks to Exwayz software [Exwayz(2024)]. Due to the high resolution of the headlight we were able to collect data with smaller cell size than in simulation, giving more insight into pattern resolution impact. The performance of our model on this dataset demonstrate LED capabilities on complex real-world scenarios. We illustrate additional qualitative results in fig.˜S5. A video of DepthFormer (LED 0.25°) qualitative results on our real-world dataset is available in the supplementary material.
Figure S6: Zoomed in qualitative results on NSDD showcasing LED improved depth prediction compared to high-beam (HB).
J Qualitative Results on Synthetic Dataset
We illustrate in fig.˜S6 the improvements of depth estimation achieved by our method by comparing results with and without pattern. They are obtained by taking the encoder-decoder, trained on high beam (HB), and with our HD pattern (LED). Red boxes emphasize on enhanced regions. Some are zoomed in for improved visibility. LED results (right) exhibit higher precision, leading to more accurate object boundaries and shapes compared to HB (middle). Far away obstacles are better defined and less blurry (first row), vehicles and pedestrians are sharper (second rows).